OpenStudy (anonymous):

Which value of k makes (2x + 3)(kx – 5) = 2kx^2 – x – 15 true? A. –3 B. –1 C. 1 D. 3

OpenStudy (anonymous):

2kx^2-10x+3kx-15=2kx^2-x-15. But then what?

OpenStudy (cwrw238):

the coefficient of the middle term ( -1) is obtained by by + 3k - 10 = -1 solve for k

OpenStudy (anonymous):

compare the coefficients ..on both the sides..! so u see k=1

OpenStudy (cwrw238):

no thats incorrect

OpenStudy (anonymous):

@Akpellet_math @cwrw238 I'm sorry but which two terms on both sides do I use?

OpenStudy (cwrw238):

solveing 3k - 10 = -1 will give you the answer

OpenStudy (anonymous):

ok thanks!

OpenStudy (cwrw238):

the coefficients of x^2 on both sides are 2k - you cant infer anything from this

OpenStudy (anonymous):

So D.

OpenStudy (anonymous):

ohh ..yes ..i am sorry ..! was in a hurry ..yes the guy is correct .!

OpenStudy (cwrw238):

thats ok